Grade 11Magical Mathematics[Interesting Approach]

3)There are 3 mixtures when mixed in the ratio of 3:4:5 price was 240.When mixed with 6:4:9 price was 340.what's the price when mixed in 5:6:8 ration price is_________

doppalapudi sairam
15 Years agoGrade 11

1 Answer

Sher Mohammad
12 Years ago
3a+6b=5c
4a+4b=6c
5a+9b=8c
a-11b=
a=11b
69b=5c
b=69*5
c=69
a=11*69*b